# On the pro-modularity in the residually reducible case for some totally real fields

### v1

@@ 1. Introduction
More precisely, suppose that $\bar{\rho}_{0}^{\textnormal{ss}}=1\oplus\bar{\chi}$ for some character $\bar{\chi}$ unramified outside $Np$, where $N$ is a positive integer. Let $R_{\bar{\rho}_{x}}$ be the universal deformation ring of $\rho_{x}$, where $0\neq x\in H^{1}(G_{\mathbb{Q},Np},\mathbb{F}(\bar{\chi}))$ and $\rho_{x}$ is the unique reducible Galois representation determined by $x$ such that $\rho_{x}^{\textnormal{ss}}=1\oplus\bar{\chi}$. Let $R_{\bar{\rho}_{0}}^{\textnormal{pd}}$ be the universal pseudo-deformation ring of $\bar{\rho}_{0}^{\textnormal{ss}}$. Deo proves the following facts under some mild assumptions .
1) [Deo23, Corollary 3.5] The ring $R_{\bar{\rho}_{x}}$ is a local complete intersection ring of Krull dimension $4$.
2) [Deo23, Theorem 4.5] The natural surjection $R_{\bar{\rho}_{x}}\twoheadrightarrow T(N)_{\bar{\rho}_{0}}$ is an isomorphism, where $T(N)_{\bar{\rho}_{0}}$ is the localization of the big Hecke algebra $T(N)$ (of level $N$) at the maximal ideal $\mathfrak{m}_{0}$ determined by $\bar{\rho}_{0}$.
In [Deo23], a crucial assumption is the cyclicity of $H^{1}(G_{\mathbb{Q},Np},\mathbb{F}(\bar{\chi}))$. Under this assumption, Deo uses the fact that the natural map $R_{\bar{\rho}_{0}}^{\textnormal{pd}}\to R_{\bar{\rho}_{x}}$ is actually a surjection proved by Kisin [Kis09, Corollary 1.4.4] in the proof of 1), and constructs a surjection $R_{\bar{\rho}_{x}}\twoheadrightarrow T(N)_{\bar{\rho}_{0}}$ induced by $R_{\bar{\rho}_{0}}^{\textnormal{pd}}\twoheadrightarrow T(N)_{\bar{\rho}_{0}}$ by the theory of Generalized Matrix Algebra in the proof of 2). However, such a cyclicity assumption is absurd when we consider a totally real field $F$ instead of $\mathbb{Q}$ due to the global characteristic formula.
@@
Let $F$ be a totally real field. Let $p>3$ be a rational prime and $\mathbb{F}$ be a finite field of characteristic $p$. Let $\Sigma$ be a finite set of finite places of $F$ containing $\Sigma_{p}$ consisting of all the places $v\mid p$ Let $\mathcal{O}$ be the ring of integers of some local field $E$ with a uniformizer $\pi$ and residue field $\mathbb{F}$, and $\chi:\operatorname{Gal}(\bar{F}/F)\to\mathcal{O}^{\times}$ a totally odd character unramified outside $\Sigma$, and $\bar{\chi}=\chi~{}\operatorname{mod}~{}\pi$. Let $F_{\Sigma}$ be the maximal extension of $F$ unramified outside $\Sigma$ and all infinite places. Let $\epsilon,\omega_{p}$ be the $p$-adic, $\operatorname{mod}~{}p$ cyclotomic character, respectively.
For an element $0\neq x\in H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1}))$, let $\rho_{x}$ be the unique non-split reducible Galois representation determined by $x$ such that $\rho_{x}^{\textnormal{ss}}=1\oplus\bar{\chi}$. Let $R_{x}^{\textnormal{aux}}$ be the universal deformation ring of $\rho_{x}$ of auxiliary type (see Definition 3.1.1).
Theorem 1.0.1 (Proposition 3.1.5).
Keep notations and assumptions as above. Assume that $F$ is an abelian totally real field. Further assume that
$[F:\mathbb{Q}]=\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1})))=\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}))).$
Then $R_{x}^{\textnormal{aux}}$ is a local complete intersection ring of Krull dimension $1+2[F:\mathbb{Q}]$.
Now we assume that $F$ is abelian of even degree and $p$ splits completely and $\mathcal{O}$ contains enough (depending on $\Sigma$, see Proposition 3.2.1) $p$-power roots of unity. For any place $v\in\Sigma\setminus\Sigma_{p}$, assume that $p\mid\operatorname{Nm}(v)-1$. Let $R_{\textnormal{aux}}^{\textnormal{ps}}$ be the universal pseudo-deformation ring of the residual pseudo-representation $1+\bar{\chi}$ with fixed determinant $\chi$. Let $\mathbb{T}_{\xi}$ be the big $p$-adic Hecke algebra (see Definition 4.1.1), and assume that $\bar{\rho}_{0}^{\textnormal{ss}}$ determines a maximal ideal $\mathfrak{m}_{\xi}$. For a prime $\mathfrak{p}$ of $R_{x}^{\textnormal{aux}}$, we say that it is pro-modular if it comes from a prime of $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$ (see Definition 4.2.2 and Definition 4.3.1 for the explicit definitions).
Theorem 1.0.2 (Theorem 4.3.2, Theorem 4.3.7).
Keep assumptions as above. Assume that the character $\chi$ satisfies the following conditions:
@@
(1) If $[F:\mathbb{Q}]\geq 7|\Sigma\setminus\Sigma_{p}|+2$, then for any irreducible component of $R_{\textnormal{aux}}^{\textnormal{ps}}$ of dimension at least $1+2[F:\mathbb{Q}]$, it is pro-modular.
(2) If $[F:\mathbb{Q}]\geq\operatorname{max}\{1+\frac{1}{2}\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1}))),7|\Sigma\setminus\Sigma_{p}|+2\}$. Then for any non-zero cocycle $x\in H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1}))$, any prime ideal of $R_{x}^{\textnormal{aux}}$ is pro-modular.
@@
Compared to Deo’s theorem [Deo23, Theorem 4.5], our result slightly relaxes the assumption $[F:\mathbb{Q}]=\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi})))$, and assumes an inequality instead.
@@
To prove such a theorem, we follow Pan’s strategy in [Pan22], and a key tool is the ”$R=\mathbb{T}$” theorem (see Theorem 4.2.4). Here we explain some differences between our method and Pan’s.
To prove [Pan22, Theorem 7.1.1], Pan found a ”potentially nice prime”, which means that it becomes nice (see Definition 4.2.2) after a soluble base change and is much easier to be constructed, in an irreducible component of the spectrum of the universal pseudo-deformation ring containing a fixed prime $\mathfrak{p}$ corresponding to the Galois representation in his theorem. In other words, to prove the Fontaine-Mazur conjecture in Pan’s case, it is not necessary to know the full pro-modularity for the universal deformation ring for the field $F$. See [Pan22, Section 7] for more details.
To prove our pro-modularity theorem, using Theorem 4.2.4, we only need to find a nice prime in each irreducible component with large dimension in the universal pseudo-deformation ring. It can be found successfully if the degree $[F:\mathbb{Q}]$ is large enough using Corollary 2.1.7, which is the main innovation of our method.
As applications to our pro-modularity arguments, we prove the following results.
Theorem 1.0.5 (Theorem 5.1.3, Theorem 5.2.1).
Keep notations as above.
$(1)$ If $[F:\mathbb{Q}]\geq 7|\Sigma\setminus\Sigma_{p}|+2$, then $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$ is equidimensional of dimension $1+2[F:\mathbb{Q}]$.
$(2)$ For any non-zero element $x\in H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1}))$, if
then $R_{x}^{\textnormal{aux}}$ is a local complete intersection ring of Krull dimension $1+2[F:\mathbb{Q}]$.
@@ 1.1. Notations and conventions
2. Basic results
In this section, we collect some results for our main theorems.
2.1. Pseudo-representations
In this subsection, we will give some results on 2-dimensional pseudo-representations. Our main references are [SW99] and [Pan22].
@@ 2.1. Pseudo-representations
Now we give a useful result in commutative algebra.
@@
We first assume that the $i^{-1}(\mathfrak{p})$ is of height at least $d+1$. As $R$ is a CNL domain, by [LE06, Proposition 8.2.29], we know that $R$ is a Nagata ring. Then the natural injection $R\hookrightarrow B$ is finite, and thus $\operatorname{dim}R=\operatorname{dim}B$. By [LE06, Theorem 8.2.39], we know that $R$ is universally catenary, and so is $B$. Hence, we have $\operatorname{dim}B/\mathfrak{p}=\operatorname{dim}B-d$ as $B$ is local. By the incomparability ([Eis13, Corollary 4.18]), we know that $i^{-1}(\mathfrak{p})$ is of dimension at least $\operatorname{dim}B-d$. From our assumption, we have $\operatorname{dim}R\geq d+1+\operatorname{dim}B-d=\operatorname{dim}B+1$, which is impossible.
Now assume that $i^{-1}(\mathfrak{p})$ is of height $d^{\prime}<d$. As $R$ is local and catenary, it is of dimension $\operatorname{dim}R-d^{\prime}$. By going up property ([Sta23, Lemma 10.36.22]), we know that $\mathfrak{p}$ is of dimension at least $\operatorname{dim}R-d^{\prime}$. As $\operatorname{dim}R=\operatorname{dim}B$, we get a contradiction. ∎
@@
2.2. Generalized matrix algebras
In this subsection, we will briefly introduce generalized matrix algebras, and use GMA as an abbreviation. Most of the details can be found in [Deo22], [Bel19] and [BC09].
Let $R$ be a CNL ring with maximal ideal $\mathfrak{m}_{R}$ and residue field $\mathbb{F}$, and we view $R$ as a topological ring under the $\mathfrak{m}_{R}$-adic topology.
Definition 2.2.1.
We say that $A=\begin{pmatrix}R&B\\ C&R\end{pmatrix}$ a topological GMA (of type $(1,1)$) over $R$, if
1) $B$ and $C$ are topological $R$-modules,
2) an element of $A$ is of the form $\begin{pmatrix}a&b\\ c&d\end{pmatrix}$ with $a,d\in R$, $b\in B$ and $c\in C$,
3) a continuous morphism $m^{\prime}:B\otimes_{R}C\to R$ of $R$-modules is given such that for all $b_{1},b_{2}\in B$ and $c_{1},c_{2}\in C$, $m^{\prime}(b_{1}\otimes c_{1})b_{2}=m^{\prime}(b_{2}\otimes c_{1})b_{1}$ and $m^{\prime}(b_{1}\otimes c_{1})c_{2}=m^{\prime}(b_{1}\otimes c_{2})c_{1}$.
Further, $A$ is a topological $R$-algebra with the addition given by
$\begin{pmatrix}a_{1}&b_{1}\\ c_{1}&d_{1}\end{pmatrix}+\begin{pmatrix}a_{2}&b_{2}\\ c_{2}&d_{2}\end{pmatrix}=\begin{pmatrix}a_{1}+a_{2}&b_{1}+b_{2}\\ c_{1}+c_{2}&d_{1}+d_{2}\end{pmatrix},$
the multiplication given by
$\begin{pmatrix}a_{1}&b_{1}\\ c_{1}&d_{1}\end{pmatrix}\cdot\begin{pmatrix}a_{2}&b_{2}\\ c_{2}&d_{2}\end{pmatrix}=\begin{pmatrix}a_{1}a_{2}+m^{\prime}(b_{1}\otimes c_{2})c_{1}&a_{1}b_{2}+d_{2}b_{1}\\ d_{1}c_{2}+a_{2}c_{1}&d_{1}d_{2}+m^{\prime}(b_{2}\otimes c_{1})\end{pmatrix},$
and the topology given by those on $R$, $B$ and $C$.
From our definition, we can observe that for any topological CNL ring $R$, $A=M_{2}(R)$ has a unique structure of topological GMA.
Definition 2.2.2.
Let $A=\begin{pmatrix}R&B\\ C&R\end{pmatrix}$ be a topological GMA over a CNL ring $R$. We say that $A^{\prime}=\begin{pmatrix}R&B^{\prime}\\ C^{\prime}&R\end{pmatrix}$ is a sub-$R$-GMA of $A$, if $B^{\prime},C^{\prime}$ are $R$-submodules of $B,C$ respectively, and the inclusion maps keep the GMA structure.
Proposition 2.2.3.
Let $A=\begin{pmatrix}R&B\\ C&R\end{pmatrix}$ be a topological GMA over a CNL ring $R$. Let $\rho:G\to R^{\times}$ be a representation of a group $G$. Assume that there exists an element $g_{0}\in G$ such that $\rho(g_{0})$ is diagonal, with diagonal terms distinct modulo $\mathfrak{m}_{R}$. Then $R[\rho(G)]$ is a sub-$R$-GMA over $A$.
Proof.
See [Bel19, Lemma 2.4.5]. ∎
Before stating the next proposition, we first give the following setup.
Let $R$ be a CNL ring with residue field $\mathbb{F}$ and $S$ an $R$-algebra (maybe not commutative). Let $A=\begin{pmatrix}R&R_{1,2}\\ R_{2,1}&R\end{pmatrix}$ be a topological GMA over $R$ which is also a quotient of $S$. Assume that $m^{\prime}(R_{1,2}\otimes_{R}R_{2,1})\subset\mathfrak{m}_{R}$. Write $e_{1}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$ and $e_{2}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$ in $A$. Then the map $\rho_{i}:S/\mathfrak{m}_{R}S\to A/\mathfrak{m}_{R}A\to e_{i}(A/\mathfrak{m}_{R}A)e_{i}\cong M_{1}(\mathbb{F})$ defines a representation of $S/\mathfrak{m}_{R}S$ for $i=1,2$. For $i,j\in\{1,2\}$, by an extension of $\rho_{j}$ by $\rho_{i}$, we mean a representation $S/\mathfrak{m}_{R}S\to\operatorname{End}_{\mathbb{F}}(V)$ together with an exact sequence of $S/\mathfrak{m}_{R}S$-modules
$0\to\rho_{i}\to V\to\rho_{j}\to 0$
Then we know that $V$ is an $\mathbb{F}$-vector space of dimension $2$. Such an extension defines an element in the module $\operatorname{Ext}_{S/\mathfrak{m}_{R}S}^{1}(\rho_{j},\rho_{i})$.
Proposition 2.2.4.
Keep assumptions as above. For $i\neq j$, there exists a natural injective $\mathbb{F}$-map:
$\iota_{i,j}:\operatorname{Hom}_{R}(R_{i,j},\mathbb{F})\hookrightarrow\operatorname{Ext}_{S/\mathfrak{m}_{R}S}^{1}(\rho_{j},\rho_{i}).$
Proof.
It is actually [BC09, Theorem 1.5.5]. We sketch a proof here.
For an element $f\in\operatorname{Hom}_{R}(R_{i,j},\mathbb{F})$, we can define an $R$-linear map $\rho_{f}:S\to A\to M_{2}(\mathbb{F})$,
$x\mapsto\begin{pmatrix}a_{i,i}(x)&f(a_{i,j}(x))\\ 0&a_{j,j}(x)\end{pmatrix},~{}x\in S,$
where
$a_{i,j}:A/\mathfrak{m}_{R}A\to e_{i}(A/\mathfrak{m}_{R}A)e_{j},~{}x\mapsto e_{i}xe_{j}$
is the canonical projection. As $m^{\prime}(R_{1,2}\otimes_{R}R_{2,1})\subset\mathfrak{m}_{R}$, we know that $\rho_{f}$ is an $\mathbb{F}$-algebra homomorphism which defines an extension of $\rho_{j}$ by $\rho_{i}$. This defines the map $\iota_{i,j}$, and consequently, the map $\rho_{f}$ defines an element $\iota_{i,j}(f)$ in $\operatorname{Ext}_{S/\mathfrak{m}_{R}S}^{1}(\rho_{j},\rho_{i})$.
We only need to show that $\iota_{i,j}$ is actually an injection. Assume that $\iota_{i,j}(f)=0$ for some $f$. Then the extension defined by $f$ is split. From our construction, we know that the map $\rho_{f}:S\to\operatorname{End}_{\mathbb{F}}(V)$ factors through $A/\mathfrak{m}_{R}A$. Thus, the extension is also split when restricted to any subalgebra of $A/\mathfrak{m}_{R}A$. We restrict it to the subalgebra $e_{i}(A/\mathfrak{m}_{R}A)e_{j}$. Then the restricted extension is
$x\mapsto\begin{pmatrix}0&f(a_{i,j}(x))\\ 0&0\end{pmatrix},~{}x\in S.$
It is split if and only if $f=0$. This proves our result. ∎
@@ 2.2. Generalized matrix algebras
If we have a representation $\rho$ and a GMA $A$ satisfying the conditions in Proposition 2.2.3, then we can take $S$ as the group algebra $R[G]$ in Proposition 2.2.4. As $R[\rho(G)]$ is a GMA over $A$ which is also a quotient of $R[G]$, we can use Proposition 2.2.4 to study the image $\rho(G)$.

### v3

@@ 1. Introduction
More precisely, suppose that $\bar{\rho}_{0}^{\textnormal{ss}}=1\oplus\bar{\chi}$ for some character $\bar{\chi}$ unramified outside $Np$, where $N$ is a positive integer. Let $R_{\bar{\rho}_{x}}$ be the universal deformation ring of $\rho_{x}$, where $0\neq x\in H^{1}(G_{\mathbb{Q},Np},\mathbb{F}(\bar{\chi}))$ and $\rho_{x}$ is the unique reducible Galois representation determined by $x$ such that $\rho_{x}^{\textnormal{ss}}=1\oplus\bar{\chi}$. Let $R_{\bar{\rho}_{0}}^{\textnormal{pd}}$ be the universal pseudo-deformation ring of $\bar{\rho}_{0}^{\textnormal{ss}}$. Deo proves the following results.
Theorem 1.0.1.
[Deo23, Theorem A, Theorem 4.5] Let $N_{0}$ be the tame Artin conductor of $\bar{\rho}_{0}$. Suppose $N_{0}|N$, $\bar{\chi}|_{G_{\mathbb{Q}_{p}}}\neq\mathbf{1},\omega_{p}^{-1}$, $\operatorname{dim}_{\mathbb{F}}H^{1}(G_{\mathbb{Q},Np},\mathbb{F}(\bar{\chi}))=1$ and $p\nmid\phi(N)$. Then
1) There is an isomorphism $(R_{\bar{\rho}_{0}}^{\textnormal{pd}})^{\operatorname{red}}\cong T(N)_{\bar{\rho}_{0}}$, where $T(N)_{\bar{\rho}_{0}}$ is the localization of the big Hecke algebra $T(N)$ (of level $N$) at the maximal ideal $\mathfrak{m}_{0}$ determined by $\bar{\rho}_{0}$.
2) There exists an isomorphism $R_{\bar{\rho}_{x}}\twoheadrightarrow T(N)_{\bar{\rho}_{0}}$.
In his proof, a crucial assumption is the cyclicity of $H^{1}(G_{\mathbb{Q},Np},\mathbb{F}(\bar{\chi}))$. Under this assumption, Deo uses the fact that the natural map $R_{\bar{\rho}_{0}}^{\textnormal{pd}}\to R_{\bar{\rho}_{x}}$ is actually a surjection proved by Kisin [Kis09, Corollary 1.4.4] in the proof of 1), and constructs a surjection $R_{\bar{\rho}_{x}}\twoheadrightarrow T(N)_{\bar{\rho}_{0}}$ induced by $R_{\bar{\rho}_{0}}^{\textnormal{pd}}\twoheadrightarrow T(N)_{\bar{\rho}_{0}}$ by the theory of Generalized Matrix Algebra in the proof of 2). However, such a cyclicity assumption is absurd when we consider a totally real field $F$ instead of $\mathbb{Q}$ due to the global characteristic formula.
@@
Let $p>3$ be a rational prime. Now we assume that $F$ is an abelian totally real field of even degree such that $p$ splits completely in $F$. Let $\Sigma$ be a finite set of finite places of $F$ containing $\Sigma_{p}$ consisting of all the places $v\mid p$ Let $\mathcal{O}$ be the ring of integers of some local field $E$ with a uniformizer $\pi$ and residue field $\mathbb{F}$, and $\chi:\operatorname{Gal}(\bar{F}/F)\to\mathcal{O}^{\times}$ a totally odd character unramified outside $\Sigma$, and $\bar{\chi}=\chi~{}\operatorname{mod}~{}\pi$. Let $F_{\Sigma}$ be the maximal extension of $F$ unramified outside $\Sigma$ and all infinite places. Let $\epsilon,\omega_{p}$ be the $p$-adic, $\operatorname{mod}~{}p$ cyclotomic character, respectively.
For an element $0\neq x\in H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1}))$, let $\rho_{x}$ be the unique non-split reducible Galois representation determined by $x$ such that $\rho_{x}^{\textnormal{ss}}=1\oplus\bar{\chi}$. Let $R_{x}^{\textnormal{aux}}$ be the universal deformation ring of $\rho_{x}$ of auxiliary type (see Definition 3.1.1). Let $R_{\textnormal{aux}}^{\textnormal{ps}}$ be the universal pseudo-deformation ring of the residual pseudo-representation $1+\bar{\chi}$ with fixed determinant $\chi$. Let $\mathbb{T}_{\xi}$ be the big $p$-adic Hecke algebra (see Definition 4.1.1), and assume that $\bar{\rho}_{0}^{\textnormal{ss}}$ determines a maximal ideal $\mathfrak{m}_{\xi}$. For a prime $\mathfrak{p}$ of $R_{\textnormal{aux}}^{\textnormal{ps}}$ (resp. $R_{x}^{\textnormal{aux}}$), we say that it is pro-modular if it comes from a prime of $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$ (see Definition 4.2.3 and Definition 4.4.1 for the explicit definitions).
Theorem 1.0.2 (Theorem 4.4.2, Theorem 4.4.6 & Theorem 4.5.2).
Keep notations as above. For any place $v\in\Sigma\setminus\Sigma_{p}$, assume that $p\mid\operatorname{Nm}(v)-1$. Assume that the character $\chi$ satisfies the following conditions:
@@
(1) If $[F:\mathbb{Q}]\geq 7|\Sigma\setminus\Sigma_{p}|+4$, then for any irreducible component of $R_{\textnormal{aux}}^{\textnormal{ps}}$ of dimension at least $1+2[F:\mathbb{Q}]$, it is pro-modular of dimension $1+2[F:\mathbb{Q}]$.
(2) If $[F:\mathbb{Q}]\geq\operatorname{max}\{1+\frac{1}{2}\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1}))),7|\Sigma\setminus\Sigma_{p}|+4\}$. Then for any non-zero cocycle $x\in H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1}))$, any prime ideal of $R_{x}^{\textnormal{aux}}$ is pro-modular. Further, $R_{x}^{\textnormal{aux}}$ is a local complete intersection ring of Krull dimension $1+2[F:\mathbb{Q}]$.
To prove such a theorem, we follow Pan’s strategy in [Pan22], and a key tool is the patching argument (Theorem 4.2.5) at a nice prime (see Definition 4.2.3). Using Theorem 4.2.5, we only need to find a nice prime in each irreducible component of large dimension in the universal pseudo-deformation ring. It can be found successfully if the degree $[F:\mathbb{Q}]$ is large enough using Corollary 2.1.7.
@@
One may ask how we can use these pro-modularity results.
The first part of Theorem 1.0.2, i.e., the pro-modularity of the universal pseudo-deformation ring, is actually a complement of the first step in Pan’s proof of the Fontaine-Mazur conjecture in the residually reducible case (see [Pan22, page 1032]) so that after an abelian base change, the conjecture is a direct consequence of (1) of Theorem 1.0.2 and [Pan22, Corollary 3.5.10], although the full pro-modularity is not needed in his proof. In fact, Pan’s strategy is to find a “potentially nice prime”, i.e., it becomes nice after some soluble base change, and it is much easier to construct. Then he can identify some irreducible components of the universal (pseudo-)deformation ring as in the big Hecke algebra, which is enough for his purpose.
However, in the missing case of the Fontaine-Mazur conjecture in [Pan22], i.e., $p=3$ and $\bar{\chi}|_{G_{F_{v}}}=\omega_{p}^{-1}$ for any $v\mid p$, to study the pro-modularity of some irreducible component of the universal (pseudo-)deformation ring, we need to know the pro-modularity of some other ones, which means we actually need a (potentially) big $R=\mathbb{T}$ theorem in this case. More precisely, the “dihedral condition” (the third one in (2) of Definition 4.2.3) sometimes obstructs the existence of the potentially nice primes, whereas the strategy of finding nice primes in this article, especially the application of Corollary 2.1.7, still works so that we just need to find pro-modular primes rather than nice primes (see Proposition 4.5.1). In other words, our method in this paper, combining some classical patching arguments, can be able to deal with some subtle cases of the conjecture, especially the case $p=3$ and $\bar{\chi}=\omega_{3}$. Therefore, it allows us to drop Pan’s second assumption of [Pan22, Theorem 1.0.4] and hence concludes the two-dimensional Fontaine-Mazur conjecture in the regular case for odd primes. See Remark 4.5.3 and [Zha24, Section 5] for more details.
We expect that our method can be generalized to prove more modularity or automorphy lifting theorems in the residually reducible case.
@@
One may also ask whether we can find some examples such that our theorems may apply.
Our assumptions on the global character $\chi$ roughly state that it “comes from” the determinant of a continuous two-dimensional geometric Galois representation $\rho$ of $G_{\mathbb{Q}}$. In other words, starting from such a character $\chi=\operatorname{det}\rho$, using Grunwald-Wang’s theorem (see [NSW13, Theorem 9.2.8]), we can find a totally real abelian number field $F$ satisfying all of our assumptions of (1) of Theorem 1.0.2. In other words, our pro-modularity result can be viewed as a potential big $R=\mathbb{T}$ theorem for general abelian totally real fields and global characters. This is also the way of applying this theorem to prove the Fontaine-Mazur conjecture.
Remark 1.0.5.
We point out that compared to Deo’s results, our pro-modularity result on the universal pseudo-deformation ring does not need the restrictive assumption about the dimension of the cohomology group.
Our pro-modularity result on the universal deformation ring is an analogue to Deo’s result [Deo23, Theorem 4.5]. It is related to the dimension of the cohomology group since we need to bound the reducible locus of the deformation ring as in [SW99].
@@ 1.1. Notations and conventions
Acknowledgement.
2. Pseudo-representations
Pseudo-representations are the main tools to study the modularity of Galois representations in the residually reducible case. We recommend the article [BC09] to readers for a general theory of pseudo-representations.
2.1. Two-dimensional pseudo-representations
In this section, we will give some results on 2-dimensional pseudo-representations. Our main references are [SW99] and [Pan22].
@@ 2.1. Two-dimensional pseudo-representations
We first recall a useful result in commutative algebra.
@@
This is [Eis13, Proposition 9.2].
@@
This result is also used in [Zha24]. See [Zha24, Lemma 4.2.5 & 5.1.2].